OpenStudy (ambermarie151):

If two line are perpendicular and one has a slope of 1/6, what is the slope of the other line ? A.1/6 B.-1/6 C.-6 D.6

OpenStudy (lncognlto):

Do you know the formula to calculate this? :)

OpenStudy (lncognlto):

Well, m x m(perp) = -1. So you have (1/6) x m(perp) = -1.

OpenStudy (paledeer):

It would be C. To find the slope of a line that is perpendicular you use the opposite reciprocal. So you flip the fraction and change the sign

OpenStudy (ambermarie151):

I thought it was D

OpenStudy (lncognlto):

@ambermarie151 Rearranging the equation gives m(perp) = -1/(1/6) or -1(6) = -6